# SPHERE: A Hierarchical Evaluation on Spatial Perception and Reasoning for Vision-Language Models

### v1

@@ Abstract
Current vision-language models may incorporate single-dimensional spatial cues, such as depth, object boundary, and basic spatial directions (e.g. left, right, front, back), yet often lack the multi-dimensional spatial reasoning necessary for human-like understanding and real-world applications. To address this gap, we develop SPHERE (Spatial Perception and Hierarchical Evaluation of REasoning), a hierarchical evaluation framework with a new human-annotated dataset to pinpoint model strengths and weaknesses, advancing from single-skill tasks to multi-skill tasks, and ultimately to complex reasoning tasks that require the integration of multiple spatial and visual cues with logical reasoning. Benchmark evaluation of state-of-the-art open-source models reveal significant shortcomings, especially in the abilities to understand distance and proximity, to reason from both allocentric and egocentric viewpoints, and to perform complex reasoning in a physical context. This work underscores the need for more advanced approaches to spatial understanding and reasoning, paving the way for improvements in vision-language models and their alignment with human-like spatial capabilities. The dataset will be open-sourced upon publication.
SPHERE: A Hierarchical Evaluation on Spatial Perception and Reasoning for Vision-Language Models
Wenyu Zhang1, Wei En Ng2, Lixin Ma3*, Yuwen Wang2*, Jungqi Zhao4*, Boyang Li4, Lu Wang1 1Institute for Infocomm Research (IR), Agency for Science, Technology and Research (A*STAR) 2National University of Singapore (NUS) 3Tongji Unversity 4 Nanyang Technological University (NTU)
@@ 1 Introduction
Spatial perception and reasoning plays an essential role in how vision-language models (VLMs) understand and potentially interact with complex, context-rich environments Du et al. (2024); Liu et al. (2023a); Zeng et al. (2024). These capabilties are crucial for the deployment of VLMs in physical-world applications such as robotics, embodied AI, or human-assistive systems (Nair et al., 2022; Brohan et al., 2023; Stone et al., 2023; Leal et al., 2023).
@@
Current research efforts to enhance spatial sensing in VLMs have incorporated features such as depth information, object bounding boxes, and simple spatial directions (e.g. left/right, inside/outside, front/back) into the models Chen et al. (2024); Zhao et al. (2023); remyx.ai (2024a); Cai et al. (2024); Cheng et al. (2024). These state-of-the-art VLMs can leverage the information to gain better understanding of object positions and relative distances. However, the enhancements often focus on isolated, simplistic spatial cues, and therefore limit the models’ ability to handle the complexities of real-world scenarios. Practical applications demand more sophisticated spatial perception and reasoning capabilities, akin to human-like understanding. Figure 1 demonstrates failure cases of the state-of-the-art GPT-4o. In the second question, GPT-4o shows difficulty counting while interpreting object proximity. In the third question, GPT-4o understands that laptop is at the edge of the desk, but fails to reason that the laptop cannot be moved further back.
@@
In this work, we propose a new benchmarking framework called SPHERE (Spatial Perception and Hierarchical Evaluation of REeasoning). In SPHERE, we seek to systematically disentangle and measure spatial skills, in order to pinpoint model strengths and weaknesses and guide future research. We are motivated by prior research findings West et al. (2023); Tiong et al. (2024) which reveal fundamental differences between machine and human intelligence. For example, while large language models (LLMs) excel in generating coherent and contextually relevant responses for advanced topics, they often exhibit surprising errors in basic understanding, even within the content they produce West et al. (2023). We cannot assume good performance in one task to always transfer to a seemingly correlated or even more basic task.
@@
We propose SPHERE as a hierarchical evaluation framework that covers a spectrum of basic and advanced tasks. We begin with single-skill tasks (i.e. object positions, distances, sizes and counts), then expand to multi-skill tasks that require multiple spatial and visual skills (e.g. counting objects at specific positions, counting objects at certain distances, etc.). We also design two novel reasoning tasks (i.e. reasoning about occluded objects and object manipulation) that require advanced understanding of a scene as a 3-dimensional environment and the objects as physical entities that can be interacted with. We measure reasoning ability as the ability to infer about scenes or objects not explicitly observed but physically plausible in the input images. We illustrate the SPHERE framework in Figure 2. These challenging tasks allow us to assess the alignment of VLMs with human-like spatial reasoning capabilities. Our contributions are summarized as follows:
@@
We design a hierarchical evaluation framework to study the spatial perception and reasoning capabilities of models, progressing from single-skill tasks, multi-skill tasks, to complex reasoning tasks. The reasoning tasks are further decomposed into simpler sub-tasks to disentangle spatial perception and reasoning.
@@
We evaluate several state-of-the-art open-source vision-language models and reveal significant shortcomings in their capabilities. These models especially lack the ability to understand distance and proximity, to reason from both allocentric and egocentric viewpoints, and to perform complex reasoning in a physical world context.
@@ 2.1 Spatial Capabilities in Multimodal LLMs
The majority of literature on multimodal LLMs focus on their conversational ability to respond fluently to the queries of human users through natural language (Dai et al., 2023; Liu et al., 2023b, 2024; Lyu et al., 2023) or other modalities such as audio and image Wu et al. (2023); Tang et al. (2024). More recently, works have started to explore the instillation of spatial capabilities in LLMs to better understand the physical world. Zheng et al. (2024) developed BAT based on text and audio by synthesizing data with a spatial audio simulator, and Devnani et al. (2024) trained a spatially-aware audio and text embedding model. SpatialBot Cai et al. (2024) directly includes pairs of RGB and depth images in existing datasets for training. A popular approach for vision-language models is to use existing algorithms and models (e.g. object detection, segmentation, depth estimation) to extract additional information from available data, and then incorporate these enhanced data for training or finetuning (Chen et al., 2024; Zhao et al., 2023; remyx.ai, 2024a; Cai et al., 2024; Cheng et al., 2024). The resulting models have improved capabilities to infer spatial relationships between objects and to estimate object dimensions from natural images.
@@ 2.2 Benchmark Evaluations
Spatial reasoning is necessary for VLMs to understand and interact with the physical world. This has spurred the development of numerous benchmarks that vary in scope, tasks, and the types of spatial skills they focus on. EmbSpatial-Bench Du et al. (2024) evaluates only a limited number of positional relationships, namely ‘above’, ‘below’, ‘left’, ‘right’, ‘far’ and ‘close‘. Zeng et al. (2024) tests the effects of spatial relations on visual grounding tasks. VSR Liu et al. (2023a) includes proximity and topological concepts. SpatialBench Cai et al. (2024) consists a small dataset of 120 images to evaluate spatial relationships, counting and size comparison between objects. Q-Spatial Bench Liao et al. (2024) and SpatialRGPT-Bench Cheng et al. (2024) focuses on model ability to recognize quantitative information such as metric distances and object sizes. These evaluations and datasets do not disentangle the effects of multiple spatial and/or visual skills. CLEVR Johnson et al. (2016) tests compositional reasoning with multiple spatial and visual skills, but the dataset is synthetic and the objects comprise only simple 3D shapes such as spheres, cubes and cylinders. Kamath et al. (2023) isolates the effect of spatial attributes by varying the spatial relations of objects while fixing object identities, but the collected dataset focuses only on positional relationships. We study the interaction effect of a wider range of skills, and introduce a novel study of model ability to infer about scenes or objects not explicitly observed but phyiscally plausible in the input images.
@@ 3.1 Benchmark Framework
SPHERE begins at the first level with 4 single-skill tasks: Position, Counting, Distance and Size. We choose these skills because they are foundational for understanding and interacting with physical entities. At the second level, the first level skills are integrated into 3 multi-skill tasks: Position + Counting, Distance + Counting, and Distance + Size. Tasks in this category are more challenging. Specifically, Distance + Size require models to understand the concept of size constancy, instead of simply measuring size by the number of pixels occupied in an image. At the third level, 2 reasoning tasks assess models on their ability to reason about object occlusion and object manipulation. These tasks reflect scenarios commonly encountered in the physical world, such as inferring the presence or position of hidden objects or determining how objects can be moved or rearranged. Logical reasoning together with multiple spatial and visual understanding abilities may be required to successfully answer the questions in this category.
@@ 3.2 Benchmark Dataset
We manually curate and annotate a question-answering dataset for SPHERE. We use images from the test split of MS COCO-2017 Lin et al. (2014) under the Flickr Terms of Use COCO Consortium (2024) due to its rich diversity in scenes and object categories. Moreover, since COCO-2017 is a widely-used and publicly available dataset, its images are likely to be in-distribution with respect to training datasets used by VLMs, ensuring that the evaluation aligns with their learned representations. This alignment reduces potential biases stemming from dataset shifts.
Annotations for each task are conducted and cross-verified by at least two authors for consistency and accuracy. The general guidelines for annotation are:
@@ 3.2.1 Dataset Design and Statistics
We annotate a total of 2,288 question-answer pairs. The sample size for each task is delineated in Table 1. Detailed dataset statistics are provided in the Appendix A. We choose to manually annotate the dataset to ensure the quality and correctness of the annotations. As seen in Figure 1, even the state-of-the-art GPT-4o produces erroneous responses.
@@
Position + Counting: Questions require the model to count the number of specified objects that has a specified spatial relationship with another object in the image. We include questions for both egocentric and allocentric viewpoints. Example questions are shown in Figure 4.
@@
Reasoning tasks: We design 2 tasks to evaluate the model’s ability to perform logical reasoning of the 3D physical world based on its understanding of the input 2D image. In addition to the final reasoning-based questions, we include understanding-based questions as an intermediate step to check the model’s comprehension of the 2D image.

### v2

# SPHERE: Unveiling Spatial Blind Spots in Vision-Language Models Through Hierarchical Evaluation

@@ Abstract
Current vision-language models may grasp basic spatial cues and simple directions (e.g. left, right, front, back), but struggle with the multi-dimensional spatial reasoning necessary for human-like understanding and real-world applications. To address this gap, we develop SPHERE (Spatial Perception and Hierarchical Evaluation of REasoning), a hierarchical evaluation framework supported by a new human-annotated dataset. SPHERE systematically probes models across increasing levels of complexity, from fundamental skills to multi-skill integration and high-level reasoning that combines spatial, visual, and logical understanding. Benchmark evaluation of state-of-the-art models reveals significant deficiencies, especially in reasoning about distance and proximity, understanding both egocentric and allocentric perspectives, and applying spatial logic in physical contexts. These findings expose critical blind spots in existing models and underscore the need for more advanced spatial reasoning techniques, driving the development of vision-language models that align more closely with human spatial cognition. The dataset will be open-sourced upon publication.
SPHERE: Unveiling Spatial Blind Spots in
Vision-Language Models Through Hierarchical Evaluation
Wenyu Zhang1, Wei En Ng2, Lixin Ma3*, Yuwen Wang2*, Jungqi Zhao4*, Allison Koenecke5 Boyang Li4, Lu Wang1 1Institute for Infocomm Research (IR), Agency for Science, Technology and Research (A*STAR) 2National University of Singapore (NUS), 3Tongji Unversity 4 Nanyang Technological University (NTU), 5 Cornell University
@@ 1 Introduction
Spatial perception and reasoning play an essential role in how vision-language models (VLMs) understand complex, context-rich environments Du et al. (2024); Liu et al. (2023a); Zeng et al. (2024). These capabilties are crucial for the deployment of VLMs in physical-world applications such as robotics, embodied AI, or human-assistive systems (Nair et al., 2022; Brohan et al., 2023; Stone et al., 2023; Leal et al., 2023).
@@
Current research efforts to enhance spatial sensing in VLMs have incorporated features such as depth information, object bounding boxes, and simple spatial directions (e.g. left/right, inside/outside, front/back) into the models Chen et al. (2024a); Zhao et al. (2023); AI (2024a); Cai et al. (2024); Cheng et al. (2024). However, the enhancements often focus on isolated, simplistic spatial cues, and therefore limit the models’ ability to handle the complexities of real-world scenarios. Practical applications demand more sophisticated spatial perception and reasoning capabilities, akin to human-like understanding. Figure 1 demonstrates failure cases of GPT-4o. In question 2, GPT-4o has difficulty counting while interpreting object proximity. In question 3, GPT-4o understands that the laptop is at the edge of the desk, but fails to reason that the laptop cannot be moved further back.
Model Single-skill Multi-skill Reasoning Overall Human baseline 95.4 92.5 89.0 93.0 Random baseline∗ 50.0 44.3 50.0 49.1 Best model score 78.2 58.6 64.7 67.9 Rank 1 Gemini 2.0 Flash GPT-4o GPT-4o GPT-4o Rank 2 GPT-4o LLaVA-OneVision SpatialBot-RGB Gemini 2.0 Flash Rank 3 LLaVA-OneVision Gemini 2.0 Flash Llama-3.2-Vision LLaVA-OneVision
@@
In this work, we propose a new benchmarking framework called SPHERE (Spatial Perception and Hierarchical Evaluation of REeasoning), to systematically disentangle and measure spatial skills, in order to pinpoint model strengths and weaknesses and guide future research. We are motivated by prior research findings West et al. (2023); Tiong et al. (2024) which reveal fundamental differences between machine and human intelligence. While LLMs excel in generating coherent and contextually relevant responses for advanced topics, they often exhibit surprising errors in basic understanding West et al. (2023). We cannot assume good performance in one task to always transfer to a seemingly correlated or even more basic task.
@@
SPHERE, illustrated in Figure 2, is a hierarchical evaluation framework that covers a spectrum of basic and advanced tasks. We begin with single-skill tasks (i.e. object positions, distances, sizes and counts), then expand to multi-skill tasks that require multiple spatial and visual skills (e.g. counting objects at specific positions, counting objects at certain distances, etc.). We also design two novel reasoning tasks (i.e. reasoning about occluded objects and object manipulation) that require advanced understanding of a scene as a 3-dimensional environment and the objects as physical entities that can be interacted with. We measure reasoning ability as the ability to infer about scenes or objects not explicitly observed but physically plausible in the input images. These challenging tasks allow us to assess the alignment of VLMs with human-like spatial reasoning capabilities. Our contributions are summarized as follows:
@@
We design a hierarchical evaluation framework to study the spatial perception and reasoning capabilities of models, progressing from single-skill, multi-skill, to complex reasoning tasks. The reasoning tasks are further decomposed into simpler sub-tasks to disentangle spatial perception and reasoning.
@@
We evaluate several state-of-the-art VLMs and reveal significant shortcomings in their capabilities. These models especially lack the ability to understand distance and proximity, to reason from both allocentric and egocentric viewpoints, and to perform complex reasoning in a physical world context.
@@ 2.1 Spatial Capabilities in Multimodal LLMs
The majority of literature on multimodal LLMs focus on their conversational ability to respond fluently to the queries of human users through natural language (Dai et al., 2023; Liu et al., 2023b, 2024; Lyu et al., 2023) or other modalities such as audio and image Wu et al. (2023); Tang et al. (2024). More recently, works have started to explore the instillation of spatial capabilities in LLMs to better understand the physical world. Zheng et al. (2024) developed BAT based on text and audio by synthesizing data with a spatial audio simulator, and Devnani et al. (2024) trained a spatially-aware audio and text embedding model. SpatialBot Cai et al. (2024) directly includes pairs of RGB and depth images in existing datasets for training. A popular approach for vision-language models is to use existing algorithms and models (e.g. object detection, segmentation, depth estimation) to extract additional information from available data, and then incorporate these enhanced data for training or finetuning (Chen et al., 2024a; Zhao et al., 2023; AI, 2024a; Cai et al., 2024; Cheng et al., 2024). The resulting models have improved capabilities to infer spatial relationships between objects and to estimate object dimensions from natural images.
@@ 2.2 Benchmark Evaluations
Spatial reasoning is necessary for VLMs to understand and interact with the physical world. This has spurred the development of numerous benchmarks that vary in scope, tasks, and the types of spatial skills they focus on. Some benchmarks Li et al. (2024b); Yu et al. (2024) involve spatial capabilities in only a small subset of tasks. EmbSpatial-Bench Du et al. (2024) evaluates only a limited number of positional relationships, namely ‘above’, ‘below’, ‘left’, ‘right’, ‘far’ and ‘close‘. Zeng et al. (2024) tests the effects of spatial relations on visual grounding tasks. VSR Liu et al. (2023a) includes proximity and topological concepts. SpatialBench Cai et al. (2024) consists a small dataset of 120 images to evaluate spatial relationships, counting and size comparison between objects. Q-Spatial Bench Liao et al. (2024) and SpatialRGPT-Bench Cheng et al. (2024) focuses on model ability to recognize quantitative information such as metric distances and object sizes. These evaluations and datasets do not disentangle the effects of multiple spatial and/or visual skills. CLEVR Johnson et al. (2016) tests compositional reasoning with multiple spatial and visual skills, but the dataset is synthetic and the objects comprise only simple 3D shapes such as spheres, cubes and cylinders. Kamath et al. (2023) isolates the effect of spatial attributes by varying the spatial relations of objects while fixing object identities, but the collected dataset focuses only on positional relationships. We study the interaction effect of a wider range of skills, and introduce a novel study of model ability to infer about scenes or objects not explicitly observed but phyiscally plausible in the input images.
@@ 3.1 Benchmark Framework
SPHERE begins at the first level with 4 single-skill tasks: Position, Counting, Distance and Size. We choose these skills because they are foundational building blocks for higher-order tasks, and are of interest to the research community as demonstrated through existing efforts to endow VLMs with these skills Cai et al. (2024); Chen et al. (2024a); Du et al. (2024). At the second level, the first level skills are integrated into 3 multi-skill tasks: Position + Counting, Distance + Counting, and Distance + Size. We choose to combine 2 skills at a time to minimize confounding factors. Tasks in this category are more challenging. Specifically, Distance + Size require models to understand the concept of size constancy, instead of simply measuring size by the number of pixels occupied in an image. At the third level, 2 reasoning tasks assess models on their ability to reason about object occlusion and object manipulation. These tasks reflect scenarios commonly encountered in the physical world, such as inferring the presence or position of hidden objects or determining how objects can be moved or rearranged. Logical reasoning together with multiple spatial and visual understanding abilities may be required to successfully answer the questions in this category.
@@ 3.2 Benchmark Dataset
We manually curate and annotate a question-answering dataset for SPHERE, as automatic generation with VLMs remain unreliable due to accuracy limitations, as demonstrated in Table 1 and GPT-4o failure cases in Figure 1. We use images from the test split of MS COCO-2017 Lin et al. (2014) due to its rich diversity in scenes and object categories. Since COCO-2017 is a widely-used and publicly available dataset, its images are likely to be in-distribution with respect to training datasets used by VLMs, ensuring that the evaluation aligns with their learned representations. This alignment reduces potential biases stemming from dataset shifts.
Annotations for each task are conducted and cross-verified by at least two authors for consistency and accuracy. Questions that did not elicit the same responses are discarded and replaced. The general guidelines for annotation are:
@@ 3.2.1 Dataset Design and Statistics
We annotate a total of 2,285 question-answer pairs. The sample size for each task is delineated in Table 2. Detailed dataset statistics and analysis of dataset diversity are in Appendix A.
Task level Task type Sample size Single-skill Position Ego.: 172 Allo.: 185 Counting 201 Distance 202 Size 198 Multi-skill Position + Counting Ego.: 64 Allo.: 105 Distance + Counting 158 Distance + Size 199 Reasoning Object occlusion Intermed.: 202 Final: 200 Object manipulation Intermed.: 199 Final: 200
@@
Position + Counting: Questions require the model to count the number of specified objects that has a specified spatial relationship with another object in the image. We include both egocentric and allocentric viewpoints. Example questions are shown in Figure 4.
@@
Reasoning tasks: We design 2 tasks to evaluate the model’s ability to logically reason about the 3D physical world based on its understanding of the input 2D image. In addition to the final reasoning-based questions, we include understanding-based questions as an intermediate step to check the model’s comprehension of the 2D image.
@@
Model Single-skill Multi-skill Reasoning Pos. Count. Dist. Size Avg P + C D + C D + S Avg Occl. Manip. Avg Overall Human baseline 90.4 98.0 98.0 95.0 95.4 95.6 91.0 91.0 92.5 87.0 91.0 89.0 93.0 Random baseline∗ 50.0 - 50.0 50.0 50.0 - - 44.3 44.3 50.0 50.0 50.0 49.1 General purpose VLMs Phi-3.5-Vision (4B) 61.8 58.7 56.7 76.0 63.3 44.4 36.1 47.2 42.6 54.4 55.5 54.9 54.3 LLaVA-NeXT (7B) 53.6 68.7 54.5 70.9 61.9 42.0 35.4 36.1 37.8 48.7 49.1 48.9 50.6 LLaVA-OneVision (7B) 60.2 76.1 64.2 84.0 71.1 53.3 53.8 52.9 53.3 53.0 59.6 56.3 61.5 Qwen2-VL (7B) 63.2 74.1 53.0 78.6 67.2 36.1 33.5 50.8 40.1 54.0 55.2 54.6 55.0 InstructBLIP (8B) 45.6 64.7 50.4 56.2 54.2 32.0 28.5 38.1 32.8 54.8 52.7 53.7 47.0 Idefics2 (8B) 50.1 45.3 49.1 58.5 50.7 22.5 27.8 40.1 30.1 51.0 53.3 52.1 44.2 InternVL2.5 (8B) 62.2 72.6 64.6 78.3 69.4 50.2 43.8 41.8 45.3 52.7 53.6 53.1 57.3 Qwen-VL (10B) 55.9 72.9 59.2 72.0 65.0 37.2 29.9 34.7 33.9 54.3 54.3 54.3 52.0 Llama-3.2-Vision (11B) 58.4 53.5 52.9 67.2 58.0 29.8 27.7 40.8 32.8 60.7 56.1 58.4 49.7 Qwen2-VL (72B) 62.8 80.6 60.8 85.5 72.4 37.3 38.0 69.3 48.2 53.9 55.0 54.4 59.8 Llama-3.2-Vision (90B) 66.0 59.2 65.2 75.8 66.5 40.1 31.8 52.8 41.6 61.6 55.3 58.4 56.2 Gemini 2.0 Flash 68.6 82.1 75.2 86.9 78.2 50.3 52.5 49.7 50.9 48.0 49.2 48.6 61.7 GPT-4o 71.7 76.6 71.3 89.4 77.3 54.4 62.7 58.8 58.6 66.0 63.3 64.7 67.9 Spatial VLMs SpatialBot-RGB (3B) 55.9 67.7 52.8 73.0 62.3 38.5 39.2 38.7 38.8 57.6 60.3 59.0 53.6 SpatialBot (3B) 55.5 72.1 51.9 75.4 63.7 39.1 36.1 38.9 38.0 56.5 54.6 55.5 53.1 SpaceMantis (8B) 52.7 52.7 57.0 63.1 56.4 28.4 43.7 45.6 39.2 56.6 54.2 55.4 50.4 SpatialRGPT-RGB (8B) 59.3 70.1 59.2 74.6 65.8 42.6 46.2 40.9 43.2 54.5 53.2 53.8 55.3
Model Position Position + Counting Allo. Ego. Allo. Ego. General purpose VLMs Phi-3.5-Vision (4B) 77.9 44.5 53.3 29.7 LLaVA-NeXT (7B) 69.3 36.7 47.6 32.8 LLaVA-OneVision (7B) 73.7 45.7 62.9 37.5 Qwen2-VL (7B) 69.1 56.0 37.1 37.5 InstructBLIP (8B) 43.0 48.4 31.4 32.8 Idefics2 (8B) 48.5 51.9 20.0 26.6 InternVL2.5 (8B) 75.7 47.7 59.8 34.4 Qwen-VL (10B) 65.6 45.5 41.5 30.0 Llama-3.2-Vision (11B) 60.5 56.0 31.6 26.9 Qwen2-VL (72B) 70.3 55.6 36.2 35.9 Llama-3.2-Vision (90B) 72.9 58.6 41.1 38.4 Gemini 2.0 Flash 83.2 52.9 55.2 42.2 GPT-4o 79.5 63.4 62.9 40.6 Spatial VLMs SpatialBot-RGB (3B) 68.6 42.2 47.6 23.4 SpatialBot (3B) 69.7 40.2 46.7 26.6 SpaceMantis (8B) 56.2 48.8 31.4 23.4 SpatialRGPT-RGB (8B) 73.2 44.4 48.6 32.8
Model Object occlusion Object manipulation Intermed. Final Final (w/ gt intermed.) Intermed. Final Final (w/ gt intermed.) General purpose VLMs Phi-3.5-Vision (4B) 60.6 54.4 67.0 59.8 55.5 57.6 LLaVA-NeXT (7B) 60.4 48.7 54.4 58.5 49.1 48.8 LLaVA-OneVision (7B) 63.9 53.0 59.8 64.2 59.6 57.1 Qwen2-VL (7B) 63.4 54.0 60.8 55.1 55.2 55.4 InstructBLIP (8B) 48.7 54.8 57.8 52.9 52.7 45.7 Idefics2 (8B) 56.5 51.0 60.0 51.1 53.3 60.7 InternVL2.5 (8B) 63.7 52.7 58.9 62.9 53.6 57.2 Qwen-VL (10B) 57.8 54.3 57.3 54.5 54.3 49.0 Llama-3.2-Vision (11B) 52.3 60.7 62.5 53.7 56.1 57.5 Qwen2-VL (72B) 65.6 53.9 62.4 68.5 55.0 76.9 Llama-3.2-Vision (90B) 62.7 61.6 68.0 59.8 55.3 74.0 Gemini 2.0 Flash 62.4 48.0 61.0 65.5 49.2 58.3 GPT-4o 67.3 66.0 72.0 68.5 63.3 76.4 Spatial VLMs SpatialBot-RGB (3B) 59.9 57.6 65.9 60.8 60.3 54.0 SpatialBot (3B) 58.6 56.5 63.2 59.2 54.6 56.1 SpaceMantis (8B) 63.1 56.6 62.4 60.1 54.2 57.6 SpatialRGPT-RGB (8B) 61.1 54.5 63.2 60.1 53.2 58.5